# Configurable photonic simulator for quantum field dynamics

### v1

@@ Abstract
Quantum field simulators provide unique opportunities for investigating the dynamics of quantum fields through tabletop experiments. A primary drawback of standard encoding schemes is their rigidity: altering the theory, its coupling geometry, metric structure, or simulation time typically requires redesigning the experimental setup, which imposes strong constraints on the types of dynamics and theories that can be simulated. Here, we introduce the Optical Time Algorithm (OTA) as a unifying framework, enabling the efficient simulation of large classes of free quantum field dynamics using a single optical circuit design that separates the time from the Hamiltonian’s structure. By modifying the parameters of the optical elements, our method allows us to engineer timescales, coupling graphs, spacetime metrics, and boundary conditions, thereby facilitating the implementation of relativistic and non-relativistic, real- and complex-valued, short- and long-range quantum field theories on both flat and curved spacetimes. We exploit the OTA’s configurability to investigate the spreading of quantum correlations in space and time for theories with continuously varying coupling ranges. Relevant features predicted by quantum field theory can be observed on systems of $10-20$ modes, which paves the ground for experimental implementations.
@@ I Introduction
40 years after Unruh’s proposal for mimicking Hawking radiation with sound waves [1], quantum field simulators have become a central tool for probing the non-equilibrium dynamics of quantum field theories [2, 3, 4, 5, 6, 7, 8]. By virtue of their high-level of control in the quantum regime, much attention has been drawn to ultracold atoms [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], trapped-ions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and Rydberg systems [40, 41, 42], all of which allow for probing few- and many-body physics. While each platform offers distinct opportunities, common approaches come with three main limitations: they operate far from the field theory limit, are often constrained by fixed coupling geometries, and allow for coherent dynamics only on short time scales. This poses significant challenges when attempting to simulate the dynamics of entire classes of proper field theories using a single setup.
With photonic devices on the rise [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], simulating quantum fields using light is emerging as a compelling alternative [56, 57, 58, 59, 60, 61, 62]. In contrast to the majority of simulators leveraging spin degrees of freedom, photons with their bosonic nature are better suited for implementing scalar field theories – among which are theories paradigmatic in high-energy physics such as the relativistic scalar, long-range theories – being of particular relevance for quantum gravity [63] – and quantum fields in curved spacetime. This, together with the high configurability offered by integrated photonic circuits, raises the need for versatile yet efficient algorithms.
@@
In this work, we propose the Optical Time Algorithm (OTA) as a unified framework for simulating relevant free scalar field theories on photonic platforms. In contrast to existing methods [64, 65, 66, 67, 68, 69] (see [70] for an overview), our decomposition into optical elements has a clear and physically intuitive structure: while a fixed array of beam splitters followed by a single layer of squeezers implements a largely arbitrary coupling geometry, the dynamics is entirely encoded in a single layer of phase shifters. This facilitates the simulation of the non-equilibrium dynamics of various quantum field theories without the need to ever alter the experimental design.
@@
As an application of our method, we investigate information dynamics in terms of entanglement entropy and quantum mutual information [71, 72], also for non-local theories. We show that analytic predictions in the continuum for how quantum entanglement [73, 74, 75, 76] spreads in spacetime based on the quasi-particle picture [77, 78, 79, 80, 81, 82, 83] are in agreement with simulations with $10-20$ modes.
@@
The remainder of this paper is structured as follows. We set the stage by recalling the necessary tools for describing continuous-variable quantum systems in Sec. II. We present our main result, the OTA, in Sec. III, and discuss its implications (we outsourced the proof to App. A). In Sec. IV, we showcase the OTA’s versatility by detailing how five rather different types of field theories can be simulated by minimally adjusting circuit parameters. As an application, we investigate the dynamics of quantum correlations after a quench in experimentally accessible regimes in Sec. V. We discuss our findings in Sec. VI.
@@
Notation. We use natural units $\hbar=c=k_{\text{B}}=1$, denote quantum operators (classical variables) by bold (normal) letters $\bm{\rho}$ ($\gamma$), analogously for their traces $\textbf{Tr}\{\bm{\rho}\}$ ($\text{Tr}\{\gamma\}$) and work in $(1,d)$-dimensional spacetime with metric signature (+,,,)(+,-$\dots$,-)( + , - , … , - ).
@@ II Preliminaries
We begin by mapping quantum fields to optical modes and introducing fundamental concepts in quantum optics [84, 85]. In particular, we introduce Gaussian operations and optical elements – phase shifters, beam splitters, and squeezers – serving as fundamental building blocks.
@@ II.1 From quantum fields to discrete modes
Figure 1: Pictorial representation of the map between scalar quantum fields (a) and the photonic simulation (b). By tuning the optical parameters, one can realize, for instance, arbitrary coupling geometries and timescales, allowing the probing of information dynamics in spacetime in general settings.
@@ II.2 Gaussian unitaries and symplectic matrices
Free, i.e., non-interacting, field theories are characterized by second-order Hamiltonians, which result in Gaussian unitaries, as they map Gaussian states – states with Gaussian Wigner function – to Gaussian states. In this case, the most general Hamiltonian takes the form
@@
A fundamental result for our analysis is the normal mode decomposition, also referred to as Williamson’s theorem [64]. Given a $2N\times 2N$ real and positive semi-definite matrix $H$ (e.g., the Hamiltonian matrix), there exists a symplectic diagonalizing (SD) transformation $S_{\text{SD}}\in Sp(2N,\mathbb{R})$ in the sense that [86, 87]
@@
defines the orthosymplectic group. It includes all transformations that are simultaneously orthogonal and symplectic, also known as passive transformations.
@@ II.3 Optical elements
    $\begin{split}H_{\text{BS}}&=\sigma_{y}\otimes\sigma_{y},\\ S_{\text{\text{BS}}}(\theta)&=\begin{pmatrix}\cos{\theta}&\sin{\theta}&0&0\\ -\sin{\theta}&\cos{\theta}&0&0\\ 0&0&\cos{\theta}&\sin{\theta}\\ 0&0&-\sin{\theta}&\cos{\theta}\end{pmatrix},\end{split}$ (12)
    describing the interference of two modes with transmittivity $\cos^{2}(\theta)$ controlled by $\theta\in[-\pi,\pi]$ [^2] A physical beam splitter is described by $\theta\in[0,\pi/2]$. The remaining values can be realized with two additional $\pi$ phase shifts: one on the first mode before the BS, and one after the BS on mode two for $\theta\in(\pi/2,\pi]$, one before on mode two for $\theta\in[-\pi,-\pi/2)$, and one after on mode one when $\theta\in[-\pi/2,0)$..
@@
For multi-mode systems, we denote by $S_{\text{PS}}(\varphi)$ the $N$-mode phase shift with phase vector $\varphi=(\varphi_{1},\dots,\varphi_{N})^{\intercal}$ composed out of single-mode phase shifts $\varphi_{j}$; analogously for the $N$-mode squeezer. While the ordering of the individual single-mode components is irrelevant in these two cases, we fix the following triangular ordering for the $N$-mode BS (see Fig. 2)
$\begin{split}S_{\text{BS}}(\theta)&\equiv S_{\text{BS}}(\theta_{12})S_{\text{BS}}(\theta_{13})\dots S_{\text{BS}}(\theta_{1N})\\ &\hskip 39.26494pt\times S_{\text{BS}}(\theta_{23})\dots S_{\text{BS}}(\theta_{2N})\\ &\hskip 61.17325pt\times\dots\\ &\hskip 79.09862pt\times S_{\text{BS}}(\theta_{N-1N}),\end{split}$ (14)
with $S_{\text{BS}}(\theta_{jk})$ being the two-mode BS over modes $j,k$.
@@ III.1.1 Isolating time dependencies
III.1.2 OTA
For the vast majority of physically relevant scalar quantum field theories, the Hamiltonian matrix is of block-diagonal form $H=H^{\phi}\oplus H^{\pi}$ with commuting blocks $[H^{\phi},H^{\pi}]=\mathds{O}_{N}$. Under this assumption, we prove in App. A.2 that $S(t)$ admits the general decomposition
$S(t)=S_{\text{BS}}(\theta)\,S_{\text{Sq}}^{-1}(z)\,S_{\text{PS}}(\varphi(t))\,S_{\text{Sq}}(z)\,S_{\text{BS}}^{-1}(\theta).$ (16)
Remarkably, the design of the corresponding optical circuit, see Fig. 2, is independent of the theory under consideration. Instead, the structure of the Hamiltonian of interest is fully encoded in the parameters of the optical elements via
@@ III.1.2 OTA
The triangular structure of the BS layers is dictated by Eq. (14), see the zoom-in in Fig. 2. It remains to find the BS parameters, for which we apply a standard iterative scheme, see App. A.2.3. Given a Hamiltonian matrix $H$, all circuit parameters can be calculated using [90].
Figure 2: Pictorial representation of the OTA for $N=5$.
The intuition behind the OTA is straightforward: the first beam-splitter layer diagonalizes the Hamiltonian matrix, thereby decoupling the modes. The following squeezer layer equalizes the prefactors of the fields $\phi$ and $\pi$, after which the (now independent) modes are evolved in time by phase shifters. Then, the second pair of a squeezer and beam-splitter layer transforms back to the original basis. As we will discuss in Sec. IV, the phase shift parameters $d_{j}$ often reduce to the dispersion relation $\omega_{j}$ of the field theory of interest. Further, we remark that the number of gates required by the OTA scales quadratically with the number of modes $N$, i.e.,
$n_{\text{G}}=n_{\text{Sq}}+n_{\text{PS}}+n_{\text{BS}}=N(N+2),$ (20)
consistent with other unitary decompositions in terms of single- and two-mode gates [91].
III.2 OTA with a single squeezer layer
III.2.1 General case
@@
$S(t)=S_{\text{BS}}(\theta)S_{\text{PS,1}}(t)S_{\text{Sq}}(t)S_{\text{PS,2}}(t)S_{\text{BS}}^{-1}(\theta).$ (21)
@@
While the BS layers remain time-independent, both the squeezer and phase shifter arrays acquire time dependencies. This differs from the result one would obtain when directly applying the Bloch-Messiah decomposition, in which case, all gates are generally time-dependent.
What is more, the single-mode part can be summarized as so-called complex squeezers, defined by
$\begin{split}&H_{\text{CSq}}(\Phi)=-\sigma_{x}\cos\Phi-\sigma_{z}\sin\Phi,\\ &S_{\text{CSq}}(\zeta)=\cosh\xi\,\mathds{1}_{2}-\sinh\xi\begin{pmatrix}\cos\Phi&\sin\Phi\\ \sin\Phi&-\cos\Phi\end{pmatrix},\end{split}$ (22)
where $\zeta=\xi e^{i\Phi}$ with $\xi\in\mathbb{R}$ and $\Phi\in[0,\pi)$, which reduces to the standard squeezer, cf. Eq. (13), when $\Phi=0$, in which case $\zeta=\xi=z$. It is straightforward to check that
$S_{\text{Sq}}^{-1}(z)S_{\text{PS}}(\varphi)S_{\text{Sq}}(z)=S_{\text{CSq}}(\zeta),$ (23)
where the parameters on both sides are related via
$\begin{split}\cosh(2\xi)&=\cos^{2}(\varphi)+\sin^{2}(\varphi)\cosh(4z),\\ \tan\Phi&=-\dfrac{\cot\varphi}{\cosh(2z)}.\\ \end{split}$ (24)
Therefore, one can always replace the three inner single-mode layers of the OTA (16) (see also Fig. 2) with a single layer of complex squeezers. As was the case above, the price to pay is that the complex squeezer parameters $\xi$ and $\Phi$ are both time-dependent, since $\varphi_{j}(t)=d_{j}t$.
III.2.2 Coherent inputs
So far, we have worked at the level of time-evolution operators without assuming a specific input state. Interestingly, the OTA requires fewer gate layers when considering the standard resource in quantum optics experiments, that is, coherent input states. We switch to mode operators $\bm{a}_{j}=(\bm{\phi}_{j}+i\bm{\pi}_{j})/\sqrt{2}$ to define coherent states as displaced vacuum states |$\ket{\alpha}=\bm{D}(\alpha)\ket{0}$, where $\bm{D}(\alpha)=e^{\alpha\bm{a}^{\dagger}-\alpha^{*}\bm{a}}$ with $\bm{a}=(\bm{a}_{1},\dots,\bm{a}_{N})^{\intercal}$ (analogously for the vector of complex phases $\alpha\in\mathbb{C}^{N}$). Here, the vacuum state |$\ket{0}$ is the product state annihilated by all $\bm{a}_{j}$.
Noting that passive transformations map coherent states onto coherent states [84], the first BS layer merely modifies the coherent states’ amplitudes $\alpha\to\alpha^{\prime}$, i.e., the first moment of the state. Remarkably, thanks to the simplifications involving the complex squeezers in Eq. (23), the OTA can be rearranged to a layer of time-dependent squeezers and a single time-independent BS array, to wit
$S(t)\bra{\alpha}\bm{r}\ket{\alpha}=S_{\text{BS}}(\theta)S_{\text{CSq}}(\zeta(t))\bra{\alpha^{\prime}}\bm{r}\ket{\alpha^{\prime}}.$ (25)
We sketch the corresponding circuit in Fig. 3.
For the vacuum input, both coherent amplitudes vanish, i.e., |$\ket{\alpha^{\prime}}=\ket{\alpha}=\ket{0}$. In this case, the OTA reduces to processing a single-mode squeezed vacuum state input with passive linear optics, as realized in Gaussian boson samplers [92, 93, 51, 54]. Note that in the standard formulation, the Gaussian boson sampler includes particle-number-resolving detectors at the output [94, 95, 96], while no particular readout is preferred by the OTA. Instead, for quantum field simulation, the detection method can be chosen to suit the application of interest.

### v2

@@ Abstract
Quantum field simulators provide unique opportunities for investigating the dynamics of quantum fields through tabletop experiments. A primary drawback of standard encoding schemes is their rigidity: altering the theory, its coupling geometry, metric structure, or simulation time typically requires redesigning the experimental setup, which imposes strong constraints on the types of dynamics and theories that can be simulated. Here, we introduce the Optical Time Algorithm (OTA) as a unifying framework, enabling the efficient simulation of large classes of free quantum field dynamics using a single optical circuit design that separates the time from the Hamiltonian’s structure. By modifying the parameters of the optical elements, our method allows us to engineer timescales, coupling graphs, spacetime metrics, and boundary conditions, thereby facilitating the implementation of relativistic and non-relativistic, real- and complex-valued, short- and long-range quantum field theories on both flat and curved spacetimes. We exploit the OTA’s configurability to investigate the spreading of quantum correlations in space and time for theories with continuously varying coupling ranges. Relevant features predicted by quantum field theory can be observed on systems of $10$ to $20$ modes, which paves the ground for experimental implementations.
@@ I Introduction
Over 40 years after Unruh’s proposal for mimicking Hawking radiation with sound waves [150] and Feynman’s insight that quantum physics is best simulated by quantum computers [58], quantum field simulators have become a central tool for probing the non-equilibrium dynamics of quantum field theories [64, 7, 17, 73]. By virtue of their high-level of control in the quantum regime, much attention has been drawn to ultracold atoms [24, 118, 153, 145], trapped-ions [22, 31, 104], and Rydberg systems [30, 26, 68], all of which allow for probing few- and many-body physics. While each platform offers distinct opportunities, common approaches share three main limitations: they operate far from the field-theory limit, are often constrained by fixed-coupling geometries, and allow coherent dynamics only on short time scales. This poses significant challenges for simulating the dynamics of entire classes of proper field theories within a single setup.
Also photonic systems have long been considered for quantum simulation [12, 112, 137] and computing [90, 91, 128] tasks. In recent years, the community has put a strong focus on quantum-advantage tests based on boson sampling, using either Fock-state inputs [2] or Gaussian inputs [74, 92]. Remarkably, Gaussian boson sampling operates entirely in non-interacting, i.e., Gaussian, regimes until photon-number detection at the output—yet, sampling from the outcome distribution is widely believed to pose a #\#P-hard problem in general [75]. The strongest experimental indications for quantum advantage have so far emerged for the Gaussian variant, with Jiuzhang and successive upgrades [165, 164, 54, 96] as well as Borealis [98]. These demonstrations highlight what current photonic hardware excels at: large mode numbers and Hilbert-space dimensions ($\sim 10^{2}$$10^{4}$ modes [96]), highly programmable Gaussian unitaries (reaching $\sim 95\%$ similarity between measured and target amplitude distributions [140]), and fast sampling rates (up to $\sim 10^{6}$ samples/s [140]). At the same time, they emphasize photonics’ main bottleneck: loss is ubiquitous, and sufficiently lossy devices can bring medium-sized instances back within reach of efficient classical algorithms [113, 114].
Given the unique capabilities of photonic platforms, especially when integrated into programmable devices [117, 156, 5, 123], simulating quantum fields with light has emerged as a compelling alternative to current approaches. In contrast to simulators based on spin degrees of freedom, photons with their bosonic nature are better suited for implementing scalar field theories, thus overcoming the need for digitization [83, 84, 89]. Further, leveraging the programmability of photonic circuits enables simulating quantum fields in curved spacetimes beyond the constraints of Unruh’s analog-gravity paradigm [15, 29], which relies on fluid-like analogies in implementations with ultracold atoms [55, 153, 142, 106, 79], optical fibers [120], or quantum fluids [141, 143].
[…]
@@
In this work, we propose the Optical Time Algorithm (OTA) as a unified framework for simulating free scalar field theories on photonic platforms. In contrast to existing methods, our decomposition into optical elements has a clear and physically intuitive structure: while a fixed interferometer array followed by a single layer of squeezers implements a largely arbitrary coupling geometry, the dynamics is isolated in a single layer of phase shifters. This facilitates the simulation of the non-equilibrium dynamics of various quantum field theories without the need to ever alter the experimental design.
Remarkably, we find that, for initial product coherent states, the OTA reduces to the Gaussian boson sampler setup. Therefore, our algorithm not only opens programmable quantum field simulation as a novel use case for state-of-the-art sampling experiments, but also provides a pathway to practical quantum advantage, given the classical hardness of assessing the full particle-number distribution.
@@
As an application of our method, we investigate information dynamics in terms of entanglement entropy and quantum mutual information [122, 4], also for non-local theories. We show that analytic predictions in the continuum for how quantum entanglement [8, 56] spreads in spacetime based on the quasi-particle picture [34, 38, 103] are in agreement with simulations with 102010-20 modes and remain robust to photon loss and noise.
@@
The remainder of this paper is structured as follows. We set the stage by recalling the necessary tools for describing continuous-variable quantum systems in Sec. II. We present our main result, the OTA [see Eq. (15)], in Sec. III, and discuss its relation to Gaussian boson sampling. In Sec. IV, we showcase the OTA’s versatility by detailing how five rather different types of field theories can be simulated by minimally adjusting circuit parameters. As an application, we investigate the dynamics of quantum correlations after a quench in experimentally accessible regimes in Sec. V, followed by an extensive analysis of experimental imperfections in Sec. VI. We discuss our findings in Sec. VII.
@@
Notation. We use natural units $\hbar=c=k_{\text{B}}=G=1$, denote quantum operators (classical variables) by bold (normal) letters $\bm{\rho}$ ($\gamma$), analogously for their traces $\textbf{Tr}\{\bm{\rho}\}$ ($\text{Tr}\{\gamma\}$) and work in $(1,d)$-dimensional spacetime with metric signature $(+,-,\dots,-)$.
@@ II Preliminaries
We begin by mapping quantum fields to optical modes and introducing fundamental concepts in quantum optics [159, 134]. In particular, we introduce Gaussian operations and optical elements—phase shifters, beam splitters, and squeezers—serving as fundamental building blocks.
@@ II.1 From quantum fields to discrete modes
Figure 1: Pictorial representation of the map between scalar quantum fields (a) and their photonic simulation (b). By tuning the optical parameters, one can realize arbitrary coupling geometries and timescales, thereby enabling the probing of information dynamics in spacetime in general settings.
@@ II.2 Gaussian unitaries and symplectic matrices
Free, i.e., non-interacting, field theories are characterized by second-order Hamiltonians, which result in Gaussian unitaries, as they map Gaussian states—states with a Gaussian Wigner function—to Gaussian states. In this case, the most general Hamiltonian takes the form
@@
A fundamental result for our analysis is the normal-mode decomposition, also known as Williamson’s theorem [162]. Given a $2N\times 2N$ real and positive semi-definite matrix $H$ (e.g., the Hamiltonian matrix), there exists a symplectic diagonalizing (SD) transformation $S_{\text{SD}}\in Sp(2N,\mathbb{R})$ in the sense that [139, 110]
@@
defines the orthosymplectic group. It includes all transformations that are simultaneously orthogonal and symplectic, i.e., passive transformations.
@@ II.3 Optical elements
[…]
    describing the interference of two modes with transmissivity $\cos^{2}(\theta)$ controlled by $\theta\in[-\pi,\pi]$ [^2] A physical beam splitter is described by $\theta\in[0,\pi/2]$. The remaining values can be realized with two additional $\pi$ phase shifts: one on the first mode before the BS, and one after the BS on mode two for $\theta\in(\pi/2,\pi]$, one before on mode two for $\theta\in[-\pi,-\pi/2)$, and one after on mode one when $\theta\in[-\pi/2,0)$..
@@
For multi-mode systems, we denote by $S_{\text{PS}}(\varphi)$ the $N$-mode phase shift with phase vector $\varphi=(\varphi_{1},\dots,\varphi_{N})^{\intercal}$ composed out of single-mode phase shifts $\varphi_{j}$; analogously for the $N$-mode squeezer.
@@ III.1.1 Isolating time dependencies
III.1.2 Decomposition into optical elements
For the vast majority of physically relevant scalar quantum field theories, the Hamiltonian matrix is of block-diagonal form $H=H^{\phi}\oplus H^{\pi}$ with commuting blocks $H^{\phi},H^{\pi}]=\mathds{O}_{N}$. Under this assumption, we prove in App. A.2 that $S(t)$ admits the general decomposition
$S(t)=S_{\text{I}}\,S_{\text{Sq}}^{-1}(z)\,S_{\text{PS}}(\varphi(t))\,S_{\text{Sq}}(z)\,S_{\text{I}}^{-1},$ (15)
where $S_{\text{I}}$ is a time-independent passive transformation, here denoted as interferometer. Remarkably, the design of the corresponding optical circuit, see Fig. 2, is independent of the theory under consideration. Instead, the structure of the Hamiltonian of interest is fully encoded in the parameters of the optical elements via
@@ III.1.2 Decomposition into optical elements
In App. A.2.3, we consider one of the possible schemes to further decompose the interferometer layers with beam splitters only. There, we also show how gauge redundancies can be used to minimize the number of gates.
Given a Hamiltonian matrix $H$, all circuit parameters can be calculated using a dedicated software package we made publicly available under [47].
Figure 2: Pictorial representation of the OTA for $N=5$. The first and last layers can be further decomposed via standard techniques, such as Reck’s [125] and Clements’ decompositions [45]. We provide an explicit example for the former in App. A.2.3.
The intuition behind the OTA is straightforward: the first interferometer layer diagonalizes the Hamiltonian matrix, thereby decoupling the modes. The following squeezer layer equalizes the prefactors of the fields $\phi$ and $\pi$, after which the (now independent) modes are evolved in time by phase shifters. Then, the second pair of a squeezer and interferometer layer transforms back to the original basis. As we will discuss in Sec. IV, the phase shift parameters $d_{j}$ often reduce to the dispersion relation $\omega_{j}$ of the field theory of interest.
Further, we remark that the number of gates required by the OTA scales quadratically with the number of modes $N$, i.e.,
$n_{\text{G}}=n_{\text{Sq}}+n_{\text{PS}}+n_{\text{I}}\sim N(N+2),$ (19)
consistent with other unitary decompositions in terms of single- and two-mode gates [111], since each interferometer layer requires $\sim N(N-1)/2$ gates for the decomposition. However, in strong contrast to other decompositions, only the phase shifts have to be tuned to simulate dynamical evolution, while all other components remain time-independent.
III.1.3 Time-dependent Hamiltonians
When the Hamiltonian carries explicit time dependence $\bm{H}=\bm{H}(t)$, the time evolution operator in Eq. (6) becomes
$\bm{U}(t)=\mathcal{T}\exp\bigg[-i\int_{0}^{t}\mathrm{d}t^{\prime}\bm{H}(t^{\prime})\bigg],$ (20)
where $\mathcal{T}$ denotes the time-ordering operator. In general, the latter prevents isolating the time in a single phase-shift layer as in Eq. (14) since the symplectic diagonalization becomes time-dependent as well $S_{\text{SD}}=S_{\text{SD}}(t)$.
Remarkably, there are relevant cases in which the general form of the OTA, Eq. (15), remains intact. When $\bm{H}(t_{1}),\bm{H}(t_{2})]=0\,\,\forall t_{1},t_{2}$, which amounts to time-independent eigenvectors $P\neq P(t)$, time-ordering simplifies to the identity $\mathcal{T}=\mathds{1}$. This applies, for instance, to circulant Hamiltonian matrices, which are diagonalizable via a time-independent discrete Fourier transformation, see Sec. IV.1. Consequently, the interferometer remains time-independent $S_{\text{I}}\neq S_{\text{I}}(t)$, while the phases and squeezing parameters generalize to
$\varphi(t)=(\tilde{d}_{1}(t),\ldots,\tilde{d}_{N}(t))^{\intercal},$ (21a)
$z(t)=-(\ln\tilde{\gamma}_{1}(t),\ldots,\ln\tilde{\gamma}_{N}(t))^{\intercal}/2,$ (21b)
respectively, see App. A.3 for a proof. Here, $\tilde{d}_{j}$ and $\tilde{z}_{j}$ encode the now time-dependent eigenvalues
$\tilde{\lambda}_{j}^{\phi(\pi)}(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\lambda_{j}^{\phi(\pi)}(t^{\prime}).$ (22)
via Eqs. (17) and Eqs. (18), respectively. Importantly, only the single-mode optical elements carry time dependence.
III.2 Connection to Gaussian boson sampling
III.2.1 OTA with a single squeezer layer
@@
$S(t)=S_{\text{I}}S_{\text{PS,1}}(t)S_{\text{Sq}}(t)S_{\text{PS,2}}(t)S_{\text{I}}^{-1}.$ (23)
@@
While the interferometer layers remain time-independent, both the squeezer and phase shifter arrays acquire time dependencies. This differs from the result obtained by directly applying the Bloch-Messiah decomposition, in which case all gates are generally time-dependent. In fact, well-established Bloch-Messiah decomposition algorithms do not rely on standard or symplectic diagonalization to isolate the time parameter [78].
III.2.2 Vacuum inputs
[…]
where $\zeta=\xi e^{i\Phi}$ encodes amplitude $\xi\in\mathbb{R}$ and phase $\Phi\in[0,\pi)$. The latter reduces to a real squeezer, see Eq. (13), when $\Phi=0$, in which case $\zeta=\xi=z$. The optical parameters in Eqs. (16a) and (16b) are related with the ones of the complex squeezers via
$\begin{split}\cosh(2\xi)&=\cos^{2}(\varphi)+\sin^{2}(\varphi)\cosh(4z),\\ \tan\Phi&=\dfrac{\cot\varphi}{\cosh(2z)},\end{split}$ (25)
see App. A.4 for details. The price to pay is that the complex squeezer parameters become time-dependent, i.e., $\xi=\xi(t),\Phi=\Phi(t)$, since they all depend on the phases $\varphi(t)$. When considering time-dependent Hamiltonians of the form discussed in Sec. III.1.3, this structure is preserved without any additional time dependencies.